# CCSBench: Evaluating Compositional Controllability in LLMs for Scientific Document Summarization

### v1

@@ 5 Experiments
Explicit vs. Implicit Control
As shown in Table 3, implicit control proves more challenging than explicit control in the zero-shot setting.
Most models struggle with implicit tasks. For readability, the $\delta_{FKGL}$ score is less than 1 across most models, indicating limited ability to distinguish between readability levels. While GPT-4 shows a $\delta_{F}$ of 2.13, its high readability FKGL score averages nearly 17, reflecting a reading level far above what is desirable. This suggests GPT-4 falls short in enhancing readability as intended. For focus control, performance remains low, with most models achieving an F1 score around 0.5. Even GPT-4, the best performer, only reaches an F1 score of 0.71, indicating limited capacity to adapt summaries for different focus levels.
In contrast, closed-source models excel at explicit control. GPT-3.5 and GPT-4 demonstrate strong length control, with PCC scores of 0.62 and 0.84, respectively, showing a high correlation between generated and target lengths. Their MAD scores of 0.66 and 0.39 indicate minimal error in length control. Although open-source models perform weaker in this regard, all models – except for Flan-T5 – achieve high success in keyword control. Closed-source models achieve over 95% Success Rate in keyword inclusion, demonstrating strong controllability over explicit attributes.
@@
Fine-tuning with LoRA improves performance across all models. However, decoder-only (DO) models only yield minor improvements, reflecting only slight gains in controllability. For instance, the $\delta_{FKGL}$ values for settings B3 and B4 remain below 1, indicating an insufficient distinction in readability. In terms of length control, both models achieve PCC values below 0.3, reflecting a weak correlation between the generated length and the target.
In contrast, encoder-decoder (ED) models B1 and B2 achieve results comparable to or even better than GPT-4, despite their weaker performance in the zero-shot setting. B2, in particular, demonstrates strong control over length and keywords, while also showing significant improvements in readability and focus, the two implicit attributes. Under high readability requirements, the average FKGL of the generated summaries drops to below 11, corresponding to a high school reading level. For focus control, while the F1 score of 0.75 indicates potential for further improvement, it still exceeds GPT-4’s performance.
Figure 3: Average attention across all layers and heads for the last 10 tokens generated by LLaMA2 over the preceding 1,024 tokens, displayed on a log-log scale. The sharp decline in attention scores as token distance increases highlights the model’s difficulty in maintaining focus on distant tokens within the sequence.
@@
Our comparisons of B1, B2 (encoder-decoder models) with B3, B4 (decoder-only models) reveal that LLaMA2 and Mistral perform significantly worse than Flan-T5 under fine-tuning. This disparity likely stems from architectural differences. Encoder-decoder architectures, such as Flan-T5, are specifically designed for sequence-to-sequence tasks, where the encoder processes the input sequence and the decoder simultaneously generates the output by attending to the encoder’s representations. However, decoder-only models such as LLaMA2 and Mistral, handle both the source sequence and the target sequence within the same unidirectional cross-attention mechanism Fu et al. (2023). As the sequence length increases, the attention becomes increasingly diffused, which degrades the model’s ability to capture important long-term dependencies. This degradation fundamentally limits tasks requiring long-range context handling, such as controllable summarization.
To further validate this hypothesis, we analyze the attention patterns of LLaMA2 as a representative example. Figure 3 demonstrates the relationship between attention scores and token distance, revealing a sharp decline in attention as distance increases. This suggests the model focuses primarily on recent context when generating each token, progressively ignoring tokens further back in the sequence. Consequently, as sequence length increases, the model’s ability to maintain focus on control signals and relevant input diminishes, negatively impacting summarization quality.
6 Discussion
@@
Table 4: Performance comparison of Flan-T5-XXL with LoRA on single and compositional control tasks.
@@
Table 5: Case studies on conflicts between readability and focus controllability. The left column presents unchanged attribute values, while the right column shows the impact of modifying a specific attribute. Blue text highlights content that meets the conditions, while red text indicates control failures.
@@
To better understand inter-attribute conflicts, we randomly select 300 samples from the test set of $\mathcal{D}_{CC}$ for each attribute. Keeping the other attribute settings unchanged, we modified the specified attribute (e.g., changing Readability from “normal” to “high”) to create 300 new samples for each attribute accordingly. Then, we use the same model to generate the summary. For a given controllability metric, we calculate the change amplitude (CA) between the control ability before the attribute value change $m$ and after the change $m_{new}$, defined as: CA=|mnew- m m|. We illustrate the dependencies between attributes in Figure 4, where each row represents the attribute being changed and each column represents the attribute being affected. We observe that implicit attributes are more susceptible to the influence of other attributes compared to explicit attributes.
Figure 4: Attribute dependencies reflected by change amplitude (CA) on PCC (length), SR (keywords), $\delta_{FKGL}$ (readability), and F1 (focus). Each row represents the attribute being modified, and each column represents the affected attribute. Numerical values and color intensity denote the strength of the dependency.
@@
Figure 5: Average attention scores of the fusion layer for each adapter show a bias toward the focus adapter.
@@
To address these trade-offs, we explore two-stage approaches where single-attribute control is learned in the first stage, and balancing multiple attributes is fine-tuned in the second stage. Specifically, we implement LoRAHub Huang et al. (2024), which assigns a fixed weight to each LoRA module; and AdapterFusion Pfeiffer et al. (2021), which utilizes an attention layer to learn how to focus on different attribute modules. However, results in Table 3 show that neither AdapterFusion nor LoRAHub achieves meaningful controllability, with both producing poor summaries, as reflected in their low ROUGE scores. We believe this failure stems from the overly simplistic integration mechanisms, which are insufficient for learning the complexities required for compositional control, resulting in incoherent outputs.
@@
7 Conclusion
In this paper, we introduce CCSBench, a benchmark for compositional controllable summarization in the scientific domain. CCSBench includes four single-attribute datasets focused on explicit (length, keywords) and implicit (readability, empirical focus) control attributes, along with a compositional dataset that controls multiple attributes simultaneously. Our findings reveal significant limitations of popular LLMs in compositional controllability, highlighting the need for further research in this area. Additionally, we provide an in-depth analysis of the challenges models face and suggest potential directions for improvement.
Limitations
@@ 7 Conclusion
Empirical results on CCSBench reveal key challenges in LLMs’ ability to handle compositional controllable scientific summarization. While we demonstrate that fine-tuning smaller models (e.g., Flan-T5) can match GPT-4’s performance across multiple metrics, and provide deeper insights into failure causes, further research is needed to resolve our observed conflicts between control attributes. Additionally, the current version of CCSBench focuses on scientific documents written in English. We consider extending our proposed task to a broader array of languages as future work.
Appendix A Detailed Experiments Settings
A.1 Models
@@
Recent advancements in large-scale generative language models based on transformers, such as GPT Brown et al. (2020), have significantly improved performance on various natural language processing tasks. Given their widespread use and strong baseline performance, we select a range of large language models to evaluate their ability to perform compositional controlled summarization.
@@
In our experiments, we focus on a variety of model architectures, including encoder-decoder models like the Flan-T5 series, decoder-only models such as LLaMA Touvron et al. (2023) and Mistral Jiang et al. (2023), and open-source large models like GPT-4 OpenAI et al. (2024). Additionally, we evaluate their parameter-efficient fine-tuning (PEFT) versions to assess whether parameter-efficient fine-tuning enhances their ability to handle controlled summarization tasks. Moreover, we also revisit previously established baselines in the news domain for controlled summarization, namely hard prompt (HP,  Zhang et al., 2023) and soft prefix tuning (SP,  Li and Liang, 2021) approaches on smaller models. We fine-tune each of these models using our compositional control dataset, specifically designed to test their ability to handle multiple attributes simultaneously. Across these models, we explore various objectives, yielding multiple model variants tailored for controlled summarization. These configurations allow us to comprehensively study the strengths and limitations of different architectures and tuning strategies when applied to compositional summarization tasks, as described in further detail below.
@@
At test time, the model receives the input document along with a set of control signals ($doc,len,focus,kw,read$) and generates the summary accordingly. The prompt template we used can be found in the Appendix B.
@@
Following Zhang et al. (2023), in the hard prompt setting, we provide the model with explicit instructions regarding the desired control attributes by using a structured prompt format. Each control attribute is formatted as "Attribute: Value", where “Attribute” refers to one of the specified control dimensions, such as "Length", "Focus", "Keywords", or "Readability", and “Value” represents the target value for that attribute (e.g., "Noramal", "High", etc.).
@@
In this setting, we prepend external trainable parameters-referred to as "prefixes"-to each layer of the summarization model to exert fine-grained control over the generation process. For controlling each attribute value, we assign $m$ prefix embeddings for the respective attribute, where $m$ is a hyperparameter representing the length of the prefix. For instance, for controlling summary readability with a "High" specification, we assign a series of embeddings $E_{Read:high}=[e^{1}_{Read:high},...,e^{m}_{Read:high}]$, where each $e_{i}^{j}$ is a vector of the same dimensionality as the model’s word embeddings. For implementation details, readers may refer to  Li and Liang (2021).
A.2 Implementation Details
@@ A.3 Evaluation Metrics
Since focus is difficult to distinguish using simple rules, we opt to use an LLM evaluator for identification. For the generated summaries, we instruct GPT-4o to determine whether they are more conceptual-focused or empirical-focused. We experiment with various instruction patterns and validate the approach on the manually annotated dataset, achieving a best accuracy of 93%. Next, we use the categories predicted by GPT-4o as the labels for the generated summaries and calculate the F1 score for each category. A higher F1 score indicates stronger control over the focus attribute. Prompt details can be found in Appendix B.
Appendix B Prompt Templates
B.1 Readability Data Construction
# System
# Instruction
B.2 Focus Data Construction
# System
# Instruction for focus identification
Definition:
Abstract: {abstract}
# Output (Extracted by functions)
{"Empirical Content": output1, "Conceptual Content": output2}
# Instruction for paraphrasing
# Output (Extracted by functions)
{"Empirical Summary": output3, "Conceptual Summary": output4}
B.3 Focus Evaluation
# System
# Instruction
Definition:
Abstract: {abstract}
B.4 Zero-Shot LLM Experiments
# System
# Instruction
Length: The summary should fit in the specified word counts.
Document: {doc}
# Control signals
Length: 0-50 words
Keywords: "multi-scale"
Readability: high
Empirical Focus Level: high
# Output
Summary:
Appendix C Human Evaluation
@@
C.1 Instructions for Focus Evaluation
@@
Read the two scientific summaries and the referece abstract, evaluate their facuality, focus and fluency.
@@
C.2 Instructions for Readability Evaluation
Instructions to MTurk Annotators.
[Summary 1]
[Summary 2]
Questions for MTurk Annotators.
Q1: Do the two summaries agree on the main points? (Yes / No)
Q2: Which summary is more readable? (Summary 1 / Summary 2)
Q3: Is Summary 1 written in natural and fluent language? (Yes / No)
Q4: Is Summary 2 written in natural and fluent language? (Yes / No)

### v2

@@ 5. Experiments
Implicit vs. Explicit Control
As shown in Table 4, implicit control proves more challenging than explicit control.
Most models struggle with implicit tasks. For readability, the $\delta_{FKGL}$ score remains below 1 across most models, indicating poor differentiation between readability levels. While GPT-4 achieves the highest $\delta_{F}$ of 2.13, its high FKGL score ($\sim$17) suggests it generates overly complex summaries, failing to improve readability as intended. Focus control is similarly weak, with most models scoring around 0.5 F1, and even GPT-4o, the best performer, reaching only 0.72 F1, highlighting limited adaptability.
In contrast, closed-source models excel at explicit control. GPT-3.5, GPT-4, and GPT-4o achieve strong length and keyword control, with PCC scores up to 0.94, MAD scores as low as 0.09, and success rates consistently exceeding 95%. While open-source models underperform in length control, LLaMA2-7B and Mistral-7B still demonstrate strong keyword control.
Effect of Few-Shot Demonstrations
Comparing zero-shot and few-shot performance of GPT-3.5 and GPT-4, we observe that few-shot demonstrations do not improve LLM capabilities in this task. This is likely due to the challenges of processing long sequences, where the length of scientific documents limits the number of effective in-context examples. As a result, the context window becomes constrained, and overly lengthy prompts reduce model effectiveness.
@@
LoRA fine-tuning improves performance across all models, though decoder-only (DO) models show only minor gains in controllability. For example, $\delta_{FKGL}$ values for settings B3 and B4 remain below 1, indicating minimal differentiation in readability. Similarly, their length control remains weak, with PCC values below 0.3, reflecting a low correlation between generated and target lengths.
In contrast, encoder-decoder (ED) models B1 and B2 achieve results comparable to or surpassing GPT-4, despite weaker zero-shot performance. B2, in particular, exhibits strong control over length and keywords while significantly improving readability and focus—both implicit attributes. Under high readability constraints, its FKGL score drops below 11, aligning with a high school reading level. For focus control, while its F1 score of 0.75 leaves room for improvement, it still outperforms GPT-4.
Figure 3. Average attention across all layers and heads for the last 10 tokens generated by LLaMA2 over the preceding 1,024 tokens, displayed on a log-log scale. The sharp decline in attention scores as token distance increases highlights the model’s difficulty in maintaining focus on distant tokens within the sequence.
@@
Our comparison of encoder-decoder models (B1, B2) and decoder-only models (B3, B4) shows that LLaMA2 and Mistral underperform compared to Flan-T5 under fine-tuning. This disparity likely arises from architectural differences. Encoder-decoder models like Flan-T5 are optimized for sequence-to-sequence tasks, where the encoder processes input while the decoder generates output by attending to the encoded representations. In contrast, decoder-only models such as LLaMA2 and Mistral handle both source and target sequences within the same unidirectional cross-attention mechanism (Fu et al., 2023). As sequence length increases, attention becomes more diffused, weakening the model’s ability to capture long-range dependencies—critical for controllable summarization.
To validate this, we analyze LLaMA2’s attention patterns. Figure 3 shows a sharp decline in attention as token distance increases, indicating that the model prioritizes recent context while progressively disregarding earlier tokens. This degradation in long-range attention reduces the model’s ability to maintain focus on control signals and relevant input, ultimately impairing summarization quality.
6. Discussion
@@
Table 5. Performance comparison of Flan-T5-XXL with LoRA on single and compositional control tasks.
@@
Table 6. Case studies on conflicts between readability and focus controllability. The left column presents unchanged attribute values, while the right column shows the impact of modifying a specific attribute. Blue text highlights content that satisfies the control conditions, while red text indicates control failures.
@@
To better understand inter-attribute conflicts, we randomly select 300 samples from the test set of $\mathcal{D}_{CC}$ for each attribute. Keeping the other attribute settings unchanged, we modified the specified attribute (e.g., changing Readability from “normal” to “high”) to create 300 new samples for each attribute accordingly. Then, we use the same model to generate the summary. For a given controllability metric, we calculate the change amplitude (CA) between the control ability before the attribute value change $m$ and after the change $m_{new}$, defined as: CA=—mnew- m m—. We illustrate the dependencies between attributes in Figure 4, where each row represents the attribute being changed and each column represents the attribute being affected. We observe that implicit attributes are more susceptible to the influence of other attributes compared to explicit attributes.
Figure 4. Attribute dependencies reflected by change amplitude (CA) on PCC (length), SR (keywords), $\delta_{FKGL}$ (readability), and F1 (focus). Each row represents the attribute being modified, and each column represents the affected attribute. Numerical values and color intensity denote the strength of the dependency.
@@
Figure 5. Average attention scores of the fusion layer for each adapter show a bias toward the focus adapter.
@@
To address these trade-offs, we explore two-stage approaches where single-attribute control is learned in the first stage, and balancing multiple attributes is fine-tuned in the second stage. Specifically, we implement LoRAHub (Huang et al., 2024), which assigns a fixed weight to each LoRA module; and AdapterFusion (Pfeiffer et al., 2021), which utilizes an attention layer to learn how to focus on different attribute modules. However, results in Table 4 show that neither AdapterFusion nor LoRAHub achieves meaningful controllability, with both producing poor summaries, as reflected in their low ROUGE scores. We believe this failure stems from the overly simplistic integration mechanisms, which are insufficient for learning the complexities required for compositional control, resulting in incoherent outputs.
@@
Table 7. Comparison of GPT-4o’s performance on the first 200 samples of CCSBench. I+C: using introduction and conclusion sections as input; Full: using the retrieved full texts as input. Detailed descriptions of metrics are provided in Section 4.2.
Method Quality Explicit Attributes Implicit Attributes #Tokens
Length Keywords Readability Focus
[…]
Full 21.15 0.98 0.06 0.98 16.55 16.10 0.45 0.67 6.78k
I+C (ours) 20.84 0.96 0.06 0.99 15.87 15.37 0.50 0.72 1.61k
7. Discussion on Input Choice for Compositional Controllable Scientific Document Summarization
Obtaining high-quality summaries is key to effective scientific document summarization. To this end, we investigate the effects of using only the introduction and conclusion sections (I+C) versus the full scientific document as input for compositional controllable scientific document summarization.
CCSBench adopts the I+C sections as input, a choice supported by prior research demonstrating its effectiveness over full-document summarization (Cachola et al., 2020; Meng et al., 2021). This effectiveness likely stems from the challenges language models face in maintaining coherence and relevance over long contexts. As CCSBench aims to foster effective general scientific summarization rather than produce in-depth analyses, the I+C sections provide a concise yet comprehensive representation of a paper’s key elements, including task definitions, prior limitations, methodologies, and main findings (Stanford University, 2006). Thus, our CCSBench aligns with the established I+C practice in the field.
With recent LLMs such as GPT-4o (Hurst et al., 2024) supporting extended context windows of up to 128k tokens, an open question arises: Can these models overcome long-context challenges in processing full scientific documents for compositional controllable summarization? To explore this, we conduct experiments on the first 200 samples of CCSBench, retrieving their full documents and comparing GPT-4o’s summarization performance using introduction and conclusion sections (I+C) versus full-text input (Full).
The results, presented in Table 7, indicate that I+C outperforms Full in keyword preservation, readability, and focus while maintaining comparable performance on other metrics. Notably, I+C attains comparable performance while consuming only 23.7% of the tokens required for Full, validating its effectiveness and practicality for compositional controllable scientific document summarization.
8. Conclusion
We introduce CCSBench, the first benchmark for evaluating compositional controllable scientific summarization. Integrating both explicit (length, keyword inclusion) and implicit (readability, empirical focus) attributes, CCSBench provides a structured framework for assessing LLMs’ ability to generate controlled, contextually appropriate scientific summaries. Our extensive experiments establish concrete LLM baselines and reveal significant limitations in compositional controllability, emphasizing the need for further research. We identify key challenges, including trade-offs between attributes and difficulties in handling implicit control, while also outlining potential directions for improvement. By providing actionable insights into LLMs’ strengths and limitations in this challenging task, we offer a foundation for advancing the field.
Limitations and Future Work
@@ 8. Conclusion
Empirical results on CCSBench highlight key challenges in LLMs’ ability to handle compositional controllable scientific summarization. While our findings show that fine-tuning smaller models (e.g., Flan-T5) can achieve performance comparable to GPT-4 across multiple metrics and provide deeper insights into failure cases, further research is needed to better understand and resolve observed interactions between control attributes.
Currently, CCSBench focuses on scientific documents in English. Expanding this framework to support multilingual scientific summarization is a promising direction for future work. Additionally, while CCSBench is designed for effective general communication, more granular summarization – addressing the need for in-depth analysis, such as detailed experimental setups and side observations – remains an important area for exploration. Specifically, extending our framework and methodology to full scientific documents using datasets like LimGen (Faizullah et al., 2024) could provide valuable insights into more comprehensive and nuanced summarization strategies.
Acknowledgments
We are grateful to the reviewers for their constructive feedback. This work was supported by the Singapore Ministry of Education Academic Research Fund Tier 1 (251RES2216).
Appendix A Experimental Setup
A.1. Models
@@
Recent advancements in large-scale generative language models based on transformers, such as GPT (Brown et al., 2020), have significantly improved performance on various natural language processing tasks. Given their widespread use and strong baseline performance, we select a range of large language models to evaluate their ability to perform compositional controlled summarization.
@@
In our experiments, we focus on a variety of model architectures, including encoder-decoder models like the Flan-T5 series, decoder-only models such as LLaMA (Touvron et al., 2023) and Mistral (Jiang et al., 2023), and open-source large models like GPT-4 (OpenAI et al., 2024). Additionally, we evaluate their parameter-efficient fine-tuning (PEFT) versions to assess whether parameter-efficient fine-tuning enhances their ability to handle controlled summarization tasks. Moreover, we also revisit previously established baselines in the news domain for controlled summarization, namely hard prompt (HP,  Zhang et al., 2023) and soft prefix tuning (SP,  Li and Liang, 2021) approaches on smaller models. We fine-tune each of these models using our compositional control dataset, specifically designed to test their ability to handle multiple attributes simultaneously. Across these models, we explore various objectives, yielding multiple model variants tailored for controlled summarization. These configurations allow us to comprehensively study the strengths and limitations of different architectures and tuning strategies when applied to compositional summarization tasks, as described in further detail below.
@@
At test time, the model receives the input document along with a set of control signals ($doc,len,focus,kw,read$) and generates the summary accordingly. The prompt template we used can be found in the Appendix C.
@@
Following Zhang et al. (2023), in the hard prompt setting, we provide the model with explicit instructions regarding the desired control attributes by using a structured prompt format. Each control attribute is formatted as ”Attribute: Value”, where “Attribute” refers to one of the specified control dimensions, such as ”Length”, ”Focus”, ”Keywords”, or ”Readability”, and “Value” represents the target value for that attribute (e.g., ”Noramal”, ”High”, etc.).
@@
In this setting, we prepend external trainable parameters-referred to as ”prefixes”-to each layer of the summarization model to exert fine-grained control over the generation process. For controlling each attribute value, we assign $m$ prefix embeddings for the respective attribute, where $m$ is a hyperparameter representing the length of the prefix. For instance, for controlling summary readability with a ”High” specification, we assign a series of embeddings $E_{Read:high}=[e^{1}_{Read:high},...,e^{m}_{Read:high}]$, where each $e_{i}^{j}$ is a vector of the same dimensionality as the model’s word embeddings. For implementation details, readers may refer to  Li and Liang (2021).
A.2. Implementation Details
@@ A.3. Evaluation Metrics
Since focus is difficult to distinguish using simple rules, we opt to use an LLM evaluator for identification. For the generated summaries, we instruct GPT-4o to determine whether they are more conceptual-focused or empirical-focused. We experiment with various instruction patterns and validate the approach on the manually annotated dataset, achieving a best accuracy of 93%. Next, we use the categories predicted by GPT-4o as the labels for the generated summaries and calculate the F1 score for each category. A higher F1 score indicates stronger control over the focus attribute. Prompt details can be found in Appendix C.
Appendix B Human Evaluation Details
@@
B.1. Instructions for Readability Evaluation
@@
[Summary 1]
[Summary 2]
Questions for MTurk Annotators.
Q1: Do the two summaries agree on the main points? (Yes / No)
Q2: Which summary is more readable? (Summary 1 / Summary 2)
Q3: Is Summary 1 written in natural and fluent language? (Yes / No)
Q4: Is Summary 2 written in natural and fluent language? (Yes / No)
B.2. Instructions for Focus Evaluation
Instructions to MTurk Annotators.
Read the two scientific summaries and the reference abstract, evaluate their factuality, focus and fluency.
@@ B.2. Instructions for Focus Evaluation
Appendix C Prompt Templates
C.1. Readability Data Construction
# System
# Instruction
C.2. Focus Data Construction
# System
# Instruction for focus identification
Definition:
Abstract: {abstract}
# Output (Extracted by functions)
{”Empirical Content”: output1, ”Conceptual Content”: output2}
# Instruction for paraphrasing
# Output (Extracted by functions)
{”Empirical Summary”: output3, ”Conceptual Summary”: output4} ”””
C.3. Focus Evaluation
# System
# Instruction
Definition:
Abstract: {abstract}
C.4. Zero-Shot LLM Experiments
# System
# Instruction
Length: The summary should fit in the specified word counts.
Document: {doc}
# Control signals
Length: 0-50 words
Keywords: ”multi-scale”
Readability: high
Empirical Focus Level: high
# Output
Summary:
C.5. Few-Shot LLM Experiments
# System
# Instruction
Length: The summary should fit in the specified word counts.
# Examples
Examples:
Document: {doc1}
{ctrl1}
Summary: {summary1}
Document: {doc2}
{ctrl2}
Summary: {summary2}
Document: {doc3}
{ctrl3}
Summary: {summary3}
# To generate
Document: {doc}
Length: 0-50 words
Keywords: ”multi-scale”
Readability: high
Empirical Focus Level: high
# Output
Summary: